# Elastic Constants and Bending Rigidities from Long-Wavelength Perturbation Expansions

### v1

@@ I Introduction
Mechanical properties dictate almost every aspect of daily, industrial, engineering and scientific sectors. A system can be mechanically deformed by applying external loads, which originates from the microscopic nature of interatomic bonding forming it. In 1678, Hooke first obtained his empirical law to describe the force needed to linearly extend or compress a mechanical spring from its equilibrium distance [1], which states the relationship $\sigma=E\varepsilon$ in a stress-strain language with the proportional factor $E$ being the modulus of elasticity. The complete theory of linear elasticity can be elegantly described by the generalized Hooke’s law, where the constant of elastic modulus is replaced by a fourth-rank elastic tensor [2]. With this elastic tensor, macroscopic mechanical properties such as bulk, shear, and Young’s moduli, Poisson’s ratio, and Lamé’s parameters can be derived as well [3, 4, 5]. Furthermore, thermodynamic and thermomechnical properties are intimately linked with elastic tensors and moduli. One can have a rapid but crude screening of candidate materials for high-performance thermoelectrics and thermal barrier coatings using the extended Debye model and minimum thermal conductivity [6, 7, 8]. Additionally, elastic properties are also crucial in the determination of the ductile-to-brittle behavior of solids [9, 10, 11, 12, 13], elastic-energy-storage-based programmable soft machines [14], molecular crystal engineering [15, 16], and the analysis of structure and composition of planets in geophysics [17, 18, 19]. Therefore, it is of significant importance to explore mechanical and elastic properties of novel materials for a wide range of applications.
@@
Two-dimensional (2D) materials, particularly since the first exfoliation of atomically thin graphene in 2004 [20, 21], have been drawing sustained interest owing to their intriguing physical properties, such as tunable electronic band gap [22, 23], superior thermal conductivity [24, 25], ultrahigh Young’s modulus [26], quantum (anomalous) Hall effect [27, 28], and the emergence of excitonic [29, 30] and topological insulators [31]. Over the past decade, both computational design [32, 33, 34, 35] and experimental synthesis [36] have greatly enriched the family of 2D materials. Thanks to their reduced dimensionality, 2D materials can also exhibit unique and anomalous mechanical properties, including extraordinary strength [26], outstanding flexibility [37, 38], negative Poisson’s ratio [39, 40] and highly-anisotropic elastic moduli [41]. As materials become atomically thin, they can be easily corrugated by mechanical and temperature stimuli, where bending stiffness or rigidity is an essential physical parameter. For example, ripples (the thermal fluctuations of membrane height) are found ubiquitous as an intrinsic feature in ultrathin 2D crystals like graphene [42, 43, 44], which play a crucial role in the thermodynamic stability of free-standing samples [37] and the negative in-plane thermal expansion at low temperatures [45]. Besides, the out-of-plane bending properties are also of great significance for the synthesis and preparation of 2D materials [46], which eventually drives their applications in highly flexible electronic devices [47]. However, due to the challenge in performing in-situ nanomechanics to measure mechanical properties of 2D materials [48], only a few experimental data are available [49]. Especially for bending rigidities, the measurements were only performed for graphene and a few 2D transition metal dichalcogenides, including MoS2, WS2 and WSe2 [50, 51, 52]. Hence, developing precise and reliable computational tools for predicting the bending rigidities of 2D materials is highly desirable.
@@
First-principles calculations based on density functional theory (DFT) have been the workhorse to provide accurate mechanical properties of crystalline solids at relatively low cost [53, 54, 55]. There exist mainly two approaches to determine the elastic tensor of crystalline solids: finite differences [56, 57, 58, 55] and linear response method within density functional perturbation theory (DFPT) [59, 60, 61, 62, 63, 64]. In the finite-difference method, several small deformations (strains) are applied to the equilibrium lattice configuration, and the stress tensors of the resulting structures are obtained via DFT, from which the elastic constants can be calculated according to the stress-strain relation. Since elastic constants are the second-order derivatives of strain energy density with respect to strain [2, 65], an alternative approach is to compute the variation of total energy as a function of strain. The number of strains to be applied in these methods depends on the crystal system [2], where the Laue class for the centrosymmetric point groups is used to analyze the independent components of elastic tensors. While the energy-strain method requires more distorted configurations than the stress-based one, the latter is more sensitive to the convergence criteria and requires stringent computational parameters [57, 58]. By contrast, DFPT adopts the analytic differentiation of the stress under the strain perturbation in a linear-response region and solves self-consistently the Sternheimer equation [59, 60, 61, 62, 63, 64]. As a linear-response approach, the DFPT calculations generally provide more accurate elastic tensor than the finite differences whose results may depend on the magnitude of the applied strain, and it must be examined as a convergence parameter [66]. Unfortunately, the implementation of DFPT is complex and depends on the specific exchange-correlation functional and pseudopotential classes.
@@
Unlike elastic constants, bending rigidities are challenging to calculate because introducing curvature generally induces a large deformation to atomic structures and small curvatures require a large supercell with many atoms [67]. The calculations of bending properties of monolayers have thus far relied on the combination of continuum medium models with atomistic simulations using empirical interatomic potentials [68, 69, 70, 71], lacking predictive abilities. Furthermore, scaling down to nanometers, the usage of continuum medium models is also questionable due to the breakdown of the continuum approximation [71, 72]. When it comes to first-principles calculations of bending properties, a few approaches [73, 74, 75, 76, 77, 78] have been proposed to evaluate the bending rigidities of 2D materials, mostly limited to the mean curvature modulus. One method is to compute the elastic configurational energy due to curvature effect, which can be derived based on the Helfrich Hamiltonian of a membrane: $\mathcal{U}=AD^{\mathrm{P}}r^{-2}/2$ [79, 80] with $D^{\mathrm{P}}$ and $A$ being the principal (mean curvature) bending rigidity and the area of atoms per chemical formula. In this case, one needs to obtain the elastic energy $\mathcal{U}$ of nanotubular configurations at different radii $r$ and perform a fit to extract $D^{\mathrm{P}}$ [76, 81]. The Gaussian curvature modulus $D^{\mathrm{G}}$ is also accessible through this method, where a series of spheroidal fullerene-like structures with different radii are built and DFT calculations are performed to obtain the corresponding elastic energy due to Gaussian curvature: $\mathcal{U}=A(2D^{\mathrm{P}}+D^{\mathrm{G}})r^{-2}$ [76]. The main drawback of such method is the difficulty to study nanotubes with a small curvature, where each configuration contains hundreds of atoms, while the simulation with a large curvature may cause in-plane strains and related relaxation effects of atomic configuration [82]. Koskinen and Kit [73, 74, 75] introduced an approach to model the spherical membrane system using revised periodic boundary conditions (PBCs) and Bloch’s theorem which are compatible with spherical symmetry. However, it is not straightforward to adopt such revised PBCs in plane-wave DFT codes, and their applications are limited to calculations in the density-functional tight-binding formalism. Besides, cyclic DFT [77, 78] (that directly use the cyclic symmetries in the Kohn-Sham equations) has been proposed to study cylindrical systems and bending deformations. The advantage of cyclic DFT is that the calculations are carried out within one unit cell with the correct cyclic symmetries and PBCs, which is computationally efficient. The developed computational framework has been applied to study the low-curvature bending rigidities of forty-four monolayer materials [83]. Recently, Shirazian and Sauer [84] presented another first-principles approach to calculate the bending rigidities of 2D materials, by adding directly the curvature perturbation to atomic structures. They chose the bending deformed configurations that minimize membrane deformations, meanwhile preserving PBCs, and bending rigidities are then obtained by comparing the calculated bending energies to two classical models from structural mechanics. Although the mentioned ab initio approaches have been applied to study the bending deformation of 2D materials, none of them are tensorial, giving only the mean curvature bending rigidities which limits the study of elastic anisotropy.
@@
In this work, we present an efficient and accurate approach for calculating simultaneously the elastic and bending rigidity tensors of crystalline solids based on Huang’s atomic theory of elasticity [85, 86, 65], valid equivalently in any dimension. Huang [86] derived the expression for elastic stiffness tensor from a long-wavelength equation for lattice vibrations, and we extend here his method to compute also the bending rigidity tensor based on a higher-order long-wavelength perturbation expansion. Crucially, we take into account the long-range multipolar electrostatics, ensuring convergence of all the perturbative summations and the correct electric boundary conditions. We have implemented this method in the framework of modern first-principles simulations, and validate it extensively against the results available in the literature. In Sec. II, we first recall the classical continuum treatment of linear elasticity, equations of motion for elastic waves and the electromechanical coupling in dielectric medium. Then, Huang’s results on the long-wavelength acoustic vibrations and his expression for elastic stiffness tensor are derived, with the exact treatment of long-range electrostatic interactions that can be applied in the current DFT codes to correctly accounts for the electrical boundary conditions of elastic constants. In Sec. III, using a higher-order perturbation expansion, we extend the Huang’s method to describe the long-wavelength dynamics of flexural acoustic (ZA) modes in 2D materials and obtain the corresponding expression for calculating the full bending rigidity tensor. In Sec. IV, we demonstrate the accuracy and wide applicability of our method by showing example calculations on well-studied bulk and 2D materials, and compare thoroughly with both the existing experimental and theoretical results. These include silicon (Si), sodium chloride (NaCl), gallium arsenide (GaAs) and rhombohedral barium titanate (BaTiO3) for bulk solids, and monolayer graphene, molybdenum disulfide (MoS2), hexagonal boron nitride (h-BN) and indium selenide (InSe) as representative 2D materials. Finally, in Sec. V, we conclude and briefly summarize our main results in this study.
@@ II First-principles theory of elasticity
where $C_{\alpha\gamma,\beta\delta}$ is the fourth-rank elastic stiffness tensor, and $\varepsilon_{\beta\delta}$ denotes the second-rank strain tensor, with the subscripts $\alpha$, $\beta$, $\gamma$ and $\delta$ labeling the Cartesian components. Except when explicitly mentioned, we assume Einstein summation on repeated atomic positions and Cartesian directions. Apart from the induced stress field, the lattice deformation also gives rise to a change in the total energy $\mathcal{E}$ [2, 65]:
$\mathcal{E}(V,\bm{\varepsilon})-\mathcal{E}(V_{0})=V\sigma_{\alpha\gamma}\varepsilon_{\alpha\gamma}+\frac{V}{2}C_{\alpha\gamma,\beta\delta}\varepsilon_{\alpha\gamma}\varepsilon_{\beta\delta}+\mathcal{O}(\varepsilon^{3}),$ (2)
in which $V$ and $V_{0}$ are the strained and initial volumes of the system, respectively. If $V_{0}$ is the equilibrium volume, the linear term on the right-hand side of Eq. (2) vanishes, and this gives the usual definition of the strain energy density function $\mathcal{U}(\bm{\varepsilon})=\frac{1}{2}C_{\alpha\gamma,\beta\delta}\varepsilon_{\alpha\gamma}\varepsilon_{\beta\delta}$ in a quadratic form. Since the stress-strain tensors are symmetric and the elastic constant is the second-order derivative of strain energy density with respect to strain components $C_{\alpha\gamma,\beta\delta}=\partial^{2}\mathcal{U}(\bm{\varepsilon})/\partial\varepsilon_{\alpha\gamma}\partial\varepsilon_{\beta\delta}$, these conditions imply that the elastic tensor obeys the symmetry relationships
@@
The equations presented so far and the equations of motion for elastic waves introduced in the following Sec. II.1 constitute the basic continuum description of elasticity in solids.
@@ II.1 Equations of motion for elastic waves
To derive the equations of motion for homogeneous stress fields, we use the divergence theorem for stress tensor, which states that the divergence of the stress tensor is equal to the force per unit volume. Accordingly, the equations of motion for stress fields can be written as
@@
where $\rho$ is the mass density of the crystal, $t$ is the time variable, and all the spatial dependence have been omitted given the conditions of homogeneous elastic deformation. Applying the generalized Hooke’s law and substituting the strain tensor by Eq. (7) gives
@@
where $\mathbf{\overline{E}}$ and $\overline{\mathbf{P}}$ are the vectors of electric field and dielectric displacement amplitudes, respectively. The coupling between the long-wavelength elastic waves and the macroscopic electric fields shows that the interatomic forces in any type of crystals are inherently long-range, which leads to a divergent lattice summation for an infinite crystal. In particular for infrared-active solids, the long-wavelength longitudinal optical (LO) phonons have a non-analytic behavior in the vicinity of the Brillouin zone center and should be treated separately [60, 62]. In this sense, the definition of elastic tensor actually depends on the electrical boundary conditions, and a similar separate treatment for the long-range electrostatic interactions will also be needed for the calculations of elastic constants. The $C_{\alpha\gamma,\beta\delta}$ discussed so far are the “short-circuit” ones which are under the boundary condition of zero electric field and can be obtained by removing the macroscopic long-range Coulomb interactions. In contrast, the “open-circuit” elastic constants defined under the boundary condition of zero electric displacement field can be obtained by considering the additional contributions from the macroscopic electric field effects [64, 88],
@@
where the second-term is the direction-dependent piezoelectric contribution to elastic tensor with the unit vector $\hat{\mathbf{q}}$ pointing the direction of the open-circuit and zero electric displacement field. The separate treatment of the long-range multipolar interactions to obtain the elastic tensor under different electrical boundary conditions is addressed in Sec. II.4.

### v2

@@ I Introduction
Mechanical properties dictate almost every aspect of daily, industrial, engineering, and scientific sectors. A system can be mechanically deformed by applying external loads, which originates from the microscopic nature of interatomic bonding forming it. In 1678, Hooke first obtained his empirical law to describe the force needed to linearly extend or compress a mechanical spring from its equilibrium distance [1], which states the relationship $\sigma=E\varepsilon$ in a stress-strain language with the proportional factor $E$ being the modulus of elasticity. The complete theory of linear elasticity can be elegantly described by the generalized Hooke’s law, where the constant of elastic modulus is replaced by a fourth-rank elastic tensor [2]. With this elastic tensor, macroscopic mechanical properties, such as bulk, shear, and Young’s moduli, Poisson’s ratio, and Lamé’s parameters, can be derived as well [3, 4, 5]. Importantly, thermal transport and thermomechanical properties are intimately linked with elastic tensors and moduli, which enables their use in screening materials for specific applications. For example, elastic constants and the associated sound velocities are key descriptors for identifying materials with high thermal conductivity for efficient thermal management [6, 7]. One can also have a rapid but approximate screening of candidate materials for high-performance thermoelectrics and thermal barrier coatings using the extended Debye model and minimum thermal conductivity [8, 9, 10]. Furthermore, mechanical properties play a critical role in the performance of lithium-ion and solid-state batteries [11, 12]. Particularly, the design of cathode materials with high mechanical strength is essential for enhancing battery lifespan and maintaining energy capacity, as charging and discharging cycles cause dimensional changes that can lead to fracture of cathodes. In addition, elastic properties are crucial in the determination of the ductile-to-brittle behavior of solids [13, 14, 15, 16, 17], elastic-energy-storage-based programmable soft machines [18], molecular crystal engineering [19, 20], and the analysis of structure and composition of planets in geophysics [21, 22, 23]. Therefore, it is of significant importance to explore mechanical and elastic properties of novel materials for a wide range of applications.
@@
Two-dimensional (2D) materials, particularly since the first exfoliation of atomically thin graphene in 2004 [24, 25], have been drawing sustained interest owing to their intriguing physical properties, such as tunable electronic band gap [26, 27], superior thermal conductivity [28, 29], ultrahigh Young’s modulus [30], quantum (anomalous) Hall effect [31, 32], and the emergence of excitonic [33, 34] and topological insulators [35]. Over the past decade, both computational design [36, 37, 38, 39] and experimental synthesis [40] have greatly enriched the family of 2D materials. Thanks to their reduced dimensionality, 2D materials can also exhibit unique and anomalous mechanical properties, including extraordinary strength [30], outstanding flexibility [41, 42], negative Poisson’s ratio [43, 44], and highly anisotropic elastic moduli [45]. As materials become atomically thin, they can be easily corrugated by mechanical and temperature stimuli, where bending stiffness or rigidity is an essential physical parameter. For example, ripples (the thermal fluctuations of membrane height) are found ubiquitous as an intrinsic feature in ultrathin 2D crystals like graphene [46, 47, 48], which play a crucial role in the thermodynamic stability of free-standing samples [41] and the negative in-plane thermal expansion at low temperatures [49]. Besides, the out-of-plane bending properties are also of great significance for the synthesis and preparation of 2D materials [50], which eventually drives their applications in highly flexible electronic devices [51]. However, due to the challenge in performing in-situ nanomechanics to measure mechanical properties of 2D materials [52], only a few experimental data are available [53]. Especially for bending rigidities, the measurements were only performed for graphene and a few 2D transition metal dichalcogenides, including MoS2, WS2, and WSe2 [54, 55, 56]. Hence, developing precise and reliable computational tools for predicting the bending rigidities of 2D materials is highly desirable.
@@
First-principles calculations based on density functional theory (DFT) have been the workhorse to provide accurate mechanical properties of crystalline solids at relatively low cost [57, 58, 59]. There exist mainly two approaches to determine the elastic tensor of crystalline solids: finite differences [60, 61, 62, 59] and linear response method within density functional perturbation theory (DFPT) [63, 64, 65, 66, 67, 68]. In the finite-difference method, several small deformations (strains) are applied to the equilibrium lattice configuration, and the stress tensors of strained structures are obtained thanks to the introduction of the quantum theory of stress [69], from which the elastic constants can be calculated according to the stress-strain relation. Since elastic constants are the second-order derivatives of strain energy density with respect to strain [2, 70], an alternative approach is to compute the variation of total energy as a function of strain. The number of strains to be applied in these methods depends on the crystal system [2], where the Laue class for the centrosymmetric point groups is used to analyze the independent components of elastic tensors. While the energy-strain method requires more distorted configurations than the stress-based one, the latter is more sensitive to the convergence criteria and requires stringent computational parameters [61, 62]. Both finite-difference schemes require a convergence study on the magnitude of the applied strain, increasing the total number of ground-state calculations for strained systems by several times. Additionally, the structural optimization under external strains becomes computationally expensive for low-symmetry crystals and requires a high-precision calculation to identify the correct minimum [71]. By contrast, DFPT adopts the analytic differentiation of the stress under the strain perturbation in a linear-response region and solves self-consistently the Sternheimer equation [63, 64, 65, 66, 67, 68]. For a given set of DFT parameters, DFPT calculations circumvent these difficulties and generally provide numerically more accurate elastic tensors than finite differences [72]. Unfortunately, the implementation of DFPT is complex and depends on the specific exchange-correlation functional and pseudopotential classes.
@@
Unlike elastic constants, bending rigidities are challenging to calculate because introducing curvature generally induces a large deformation to atomic structures and small curvatures require a large supercell with many atoms [73]. The calculations of bending properties of monolayers have thus far relied on the combination of continuum medium models with atomistic simulations using empirical interatomic potentials [74, 75, 76, 77], lacking predictive abilities. Furthermore, scaling down to nanometers, the usage of continuum medium models is also questionable due to the breakdown of the continuum approximation [77, 78]. When it comes to first-principles calculations of bending properties, a few approaches [79, 80, 81, 82, 83, 84] have been proposed to evaluate the bending rigidities of 2D materials, mostly limited to the mean curvature modulus. One method is to compute the elastic configurational energy due to curvature effect, which can be derived based on the Helfrich Hamiltonian of a membrane: $\mathcal{U}=AD^{\mathrm{P}}r^{-2}/2$ [85, 86] with $D^{\mathrm{P}}$ and $A$ being the principal (mean curvature) bending rigidity and the area of atoms per chemical formula. In this case, one needs to obtain the elastic energy $\mathcal{U}$ of nanotubular configurations at different radii $r$ and perform a fit to extract $D^{\mathrm{P}}$ [82, 87]. The Gaussian curvature modulus $D^{\mathrm{G}}$ is also accessible through this method, where a series of spheroidal fullerene-like structures with different radii are built and DFT calculations are performed to obtain the corresponding elastic energy due to Gaussian curvature: $\mathcal{U}=A(2D^{\mathrm{P}}+D^{\mathrm{G}})r^{-2}$ [82]. The main drawback of such a method is the difficulty to study nanotubes with a small curvature, where each configuration can contain hundreds of atoms, while the simulation with a large curvature may cause prominent local distortions due to significant in-plane strains and the induced relaxation effects [88]. Koskinen and Kit [79, 80, 81] introduced an approach to model the spherical membrane system using revised periodic boundary conditions (PBCs) and Bloch’s theorem which are compatible with spherical symmetry. However, it is not straightforward to adopt such revised PBCs in plane-wave DFT codes, and their applications are limited to calculations in the density-functional tight-binding formalism. Besides, cyclic DFT [83, 84] (that directly use the cyclic symmetries in the Kohn-Sham equations) has been proposed to study cylindrical systems and bending deformations. The advantage of cyclic DFT is that the calculations are carried out within one unit cell with the correct cyclic symmetries and PBCs, which is computationally efficient. The developed computational framework has been applied to study the low-curvature bending rigidities of 44 monolayer materials [89], and a caveat is that one should verify the selected range of small curvatures yields the correct low-curvature results. Recently, Shirazian and Sauer [90] presented another first-principles approach to calculate the bending rigidities of 2D materials, by adding directly the curvature perturbation to atomic structures. They chose the bending deformed configurations that minimize membrane deformations, meanwhile preserving PBCs, and bending rigidities are then obtained by comparing the calculated bending energies to two classical models from structural mechanics. Although the mentioned ab initio approaches have been applied to study the bending deformation of 2D materials, none of them are tensorial, giving only the mean curvature bending rigidities and Gaussian modulus which limit the study of elastic anisotropy.
Here we present an efficient and accurate approach for calculating simultaneously the elastic and bending rigidity tensors of crystalline solids based on Huang’s atomic theory of elasticity [91, 92, 70], valid equivalently in any dimension. Huang [92] derived the expression for elastic stiffness tensor from a long-wavelength equation for lattice vibrations, and we extend here his method to compute also the bending rigidity tensor via a higher-order long-wavelength perturbation expansion. Crucially, we provide practical guidelines for the implementation in the framework of modern first-principles simulations, and finally demonstrated its accuracy by validating extensively against the results available in the literature. We find that good agreements can only be achieved with the subtraction of long-range multipolar electrostatics, ensuring convergence of all the perturbative summations and the correct electric boundary conditions. Therefore, we develop a novel cluster expansion approach which enables an automated, accurate, and efficient extraction of multipole and dielectric tensors up to an arbitrarily high order, based on the charge density and dielectric responses in the long-wavelength limit. Besides, our high-order extension of Huang’s theory gives access to the complete fourth-rank bending rigidity tensor, which is significant for anisotropic nanoplates, otherwise inaccessible from existing methods. This enables a direct evaluation of Gaussian curvature modulus without extra cost, and our simulations also reveal the pivotal role of lattice relaxation in 2D materials during bending deformation, establishing a fundamental platform for designing future flexible electronics.
@@
This work is organized as follows. In Sec. II, we first recall the classical continuum treatment of linear elasticity, equations of motion for elastic waves and the electromechanical coupling in dielectric medium. Then, Huang’s results on the long-wavelength acoustic vibrations and his expression for elastic stiffness tensor are derived, with the exact treatment of long-range electrostatic interactions that can be applied in the current DFT codes to correctly accounts for the electrical boundary conditions of elastic constants. In Sec. III, using a higher-order perturbation expansion, we extend the Huang’s method to describe the long-wavelength dynamics of flexural acoustic (ZA) modes in 2D materials and obtain the corresponding expression for calculating the full bending rigidity tensor. In Sec. IV, we demonstrate the accuracy and wide applicability of our method by showing example calculations on well-studied bulk and 2D materials, and compare thoroughly with both the existing experimental and theoretical results. These include silicon (Si), sodium chloride (NaCl), gallium arsenide (GaAs), and rhombohedral barium titanate (BaTiO3) for bulk solids, and monolayer graphene, molybdenum disulfide (MoS2), hexagonal boron nitride (h-BN), and indium selenide (InSe) as representative 2D materials. Finally, in Sec. V, we conclude and briefly summarize our main results in this study.
@@ II First-principles theory of elasticity
where $C_{\alpha\gamma,\beta\delta}$ is the fourth-rank elastic stiffness tensor, and $\varepsilon_{\beta\delta}$ denotes the second-rank strain tensor, with the subscripts $\alpha$, $\beta$, $\gamma$, and $\delta$ labeling the Cartesian components. Except when explicitly mentioned, we assume Einstein summation on repeated atomic positions and Cartesian directions. Apart from the induced stress field, the lattice deformation also gives rise to a change in the total energy $\mathcal{E}$ [2, 70]:
$\mathcal{E}(V,\bm{\varepsilon})-\mathcal{E}(V_{0})=\frac{V}{2}C_{\alpha\gamma,\beta\delta}\varepsilon_{\alpha\gamma}\varepsilon_{\beta\delta}+\mathcal{O}(\varepsilon^{3}),$ (2)
in which $V$ and $V_{0}$ are the strained and equilibrium volumes of the system, respectively. The linear term of Eq. (2) is not present when the system is initially at the equilibrium, and this gives the usual definition of the strain energy density function $\mathcal{U}(\bm{\varepsilon})=\frac{1}{2}C_{\alpha\gamma,\beta\delta}\varepsilon_{\alpha\gamma}\varepsilon_{\beta\delta}$ in a quadratic form. Since the stress-strain tensors are symmetric and the elastic constant is the second-order derivative of strain energy density with respect to strain components $C_{\alpha\gamma,\beta\delta}=\partial^{2}\mathcal{U}(\bm{\varepsilon})/\partial\varepsilon_{\alpha\gamma}\partial\varepsilon_{\beta\delta}$, these conditions imply that the elastic tensor obeys the symmetry relationships
@@
The equations presented so far and the equations of motion for elastic waves introduced in the following Sec. II.1 constitute the basic continuum description of elasticity in solids. While this work focuses on zero-temperature results, finite-temperature effects can be taken into account using a free energy formalism [94].
@@ II.1 Equations of motion for elastic waves
To derive the equations of motion for stress fields in the dynamic theory, we use the divergence theorem for stress tensor, which states that the divergence of the stress tensor is equal to the force per unit volume. Accordingly, the equations of motion for stress fields can be written as
@@
where $\rho$ is the mass density of the crystal, $t$ is the time variable. Applying the generalized Hooke’s law and substituting the strain tensor by Eq. (7) gives
@@
where $\mathbf{\overline{E}}$ and $\overline{\mathbf{P}}$ are the vectors of electric field and dielectric displacement amplitudes, respectively. The coupling between the long-wavelength elastic waves and the macroscopic electric fields shows that the interatomic forces in any type of crystals are inherently long-range, which leads to a divergent lattice summation for an infinite crystal. In particular for infrared-active solids, the long-wavelength longitudinal optical (LO) phonons have a non-analytic behavior in the vicinity of the Brillouin zone center and should be treated separately [64, 66]. In this sense, the definition of elastic tensor actually depends on the electrical boundary conditions, and a similar separate treatment for the long-range electrostatic interactions will also be needed for the calculations of elastic constants. The $C_{\alpha\gamma,\beta\delta}$ discussed so far are the “short-circuit” ones, which are under the boundary condition of zero electric field and can be obtained by removing the macroscopic long-range Coulomb interactions. In contrast, the “open-circuit” elastic constants defined under the boundary condition of zero electric displacement field can be obtained by considering the additional contributions from the macroscopic electric field effects [68, 95],
@@
where the second term is the direction-dependent piezoelectric contribution to elastic tensor with the unit vector $\hat{\mathbf{q}}$ pointing in the direction of the open-circuit and zero electric displacement field. The separate treatment of the long-range multipolar interactions to obtain the elastic tensor under different electrical boundary conditions is addressed in Sec. II.4.